# The Born rule – 100 years ago and today

### v1

@@ 2.3 Paradoxes and measurement context in Born’s rule
Already in 1926, Pauli [51, p.347] noticed the paradox that his objective position probability and Born’s objective momentum probability interpretation following from a spectral interpretation of the objective expectation form (BR-OE) of the Born rule are mutually incompatible. He expressed it in the following words: ”Man kann die Welt mit dem $p$-Auge und man kann sie mit dem $q$-Auge ansehen, aber wenn man beide Augen zugleich aufmachen will, dann wird man irre”
The 1927 paper by Jordan [30, p.811] cites Pauli [50] and extends Born’s second assumption further to an objective, measurement independent probability interpretation of inner products (probability amplitudes) of eigenstates of two arbitrary operators in place of position and momentum. without being aware of the conceptual problem this objective view poses when applied to noncommuting operators.
@@
2.4 Collapse
Tied to the measurement formulations of the Born rule is a controversial collapse (or state reduction) postulate about the state after a measurement. State reduction was introduced informally in Heisenberg [24, p.186] (”Jede Ortsbestimmung reduziert also das Wellenpaket wieder auf seine ursprüngliehe Grösse”) and formally established by the authority of Dirac in several editions of his book:
”The state of the system after the observation must be an eigenstate of [the observable] $\alpha$, since the result of a measurement of $\alpha$ for this state must be a certainty.” (Dirac [17, p.49], first edition, 1930)
”Thus after the first measurement has been made, the system is in an eigenstate of the dynamical variable $\xi$, the eigenvalue it belongs to being equal to the result of the first measurement. This conclusion must still hold if the second measurement is not actually made. In this way we see that a measurement always causes the system to jump into an eigenstate of the dynamical variable that is being measured, the eigenvalue this eigenstate belongs to being equal to the result of the measurement.” (Dirac [18, p.36], third edition, 1936)
Dirac’s statement was discredited (without much success) in 1958 by another authority, Landau & Lifshitz [34, Section 7], who remark that the state after the measurement is in general not an eigenstate, This is apparent in the modern notion of a quantum process (see, e.g., Chuang & Nielsen [13] or Mohseni et al. [38]), which gives the correct post-measurement state for arbitrary POVM measurements.
@@ 3 The Born rule today
In order not to follow into the same trap, it is necessary to make the place explicit where measurement enters, and to define the properties a detector must have in order that its observations count as measurements. Therefore we begin with a precise conceptual framework within which some versions of the Born rule can be derived.
Following Neumaier & Westra [43, Part II], we now give precise definitions for quantum expectations, statistical expectations, and quantum measurement devices, and derive from it the formulations (BR-POVM) and (BR-C) of the Born rule.
@@ 3.1 The detector response principle
As shown in Neumaier & Westra [43, Part II], the DRP leads naturally to all basic concepts and properties of modern quantum mechanics. In particular, it gives a precise operational meaning to quantum states, quantum detectors, quantum processes, and quantum instruments. This gives a perspective on the foundations of quantum mechanics that is quite different from the well-trodden path followed by most quantum mechanics textbooks.
@@ 3.2 The statistical interpretation of quantum mechanics
The key result for the theory of quantum measurements is the following detector response theorem, proved in [43, Section 4.1.2].
@@ 3.3 Derivation of the Born rule
It is shown in [43, Part II] that the DRP together with techniques from quantum tomography allow one to reconstruct from the DRP the complete edifice of quantum mechanics, including its dynamical and spectral consequences. In particular, the DRP is precisely what is needed to obtain the concepts and results from quantum information thoery.
@@ 3.4.4 Domain of validity of the universal form
$\bullet$ As first observed by Heisenberg [25, p.25], the Born rule implies a tiny but positive probability that an electron bound to an atom is detected light years away from the atom: ”Das Resultat ist aber merkwürdiger, als es im ersten Augenblick den Anschein hat. Bekanntlich nimmt $\psi^{*}\psi$ exponentiell mit wachsendem Abstand vom Atomkern ab. Also besteht immer noch eine endliche Wahrscheinlichkeit dafür, das Elektron in sehr weitem Abstand vom Atomkern zu finden.” Therefore $|\psi(x)|^{2}$ is unlikely to be the exact probability density for being detected at $x$, as (BR-US) would require. This indicates that for observables with continuous spectrum, (BR-US) must be an idealization.
@@ 3.4.5 Domain of validity of the condensed form
The statistical interpretation based on the Born rule in the condensed form (BR-C) and in the POVM form (BR-POVM) are derivable from the DRP, and give a correct account of the actual experimental situation. They apply without idealization to results of arbitrary quantum measurements in the sense of the DRP. For a large number of examples see Neumaier & Westra [43, Section 6.2].
However, (BR-C) and (BR-POVM) still have the same limitations as the DRP, and cannot be applied to measurement of the kind discussed at the beginning of this subsection, such as macroscopic measurements or the measurements of spectral lines. For example, spectroscopy-based high precision measurements such as the gyromagnetic ratio of the electron are not quantum measurements but measurements of numerical parameters in specific quantum models; see the detailed discussion in Neumaier & Westra [43, Section 5.2].
@@ 3.5 What is missing in the foundations?
Nature somehow produces from what has been up to (a steadily progressing) now that what has been up to a while later.
Physics is supposed to quantitatively describe and understand this in terms of physical theories and models with which predictions can be made, and in terms of experiments with which such predictions can be prompted, verified, or refuted. But Nature proceeds without caring about our physical models and the knowledge built into these, and has done so long before the first living observers existed.
The fundamental physical state $\lambda(t)$ is a complete description of the (to us unknown) collection of everything that has been up to time $t$. The $\lambda(t)$ for all $t$ fully describe what is, and the (to us unknown) fundamental dynamical law of physics determines how $\lambda(t)$ for $t>t_{0}$ is related to $\lambda(t_{0})$.
@@
Curiously, at present, our physical models are not in terms of what is (i.e., of approximations to $\lambda(t)$) but in terms of states $\psi$ or $\rho$ and observables $X$ only indirectly related to what is. With a purely epistemic quantum state, nothing real is in the theory, a state of affairs criticized by Einstein throughout his life.
@@
This indirect relation between the quantum theory of states and observables and reality is is given by one of the versions of the Born rule. These depend on a notion of measurement that has no formal counterpart in the models. This means that we are currently unable to explain theoretically which formal models of pieces of matter describe a quantum detector. The theoretical explanation for why certain objects are quantum detectors constitutes the so-called quantum measurement problem.
@@
If we accept the DRP, the problem is to give mathematical conditions on certain individual quantum objects described by microscopic quantum mechanics that guarantee that the DRP holds for a suitable mathemetcal definition of what it means for the model to respond. This is a nontrivial problem in the statistical mechanics of detector properties. The work by Allahverdyan et al. [2] goes some way towards resolving the problem. But they have to equate at some points the formal probabilities of statistical mechanics (cf. Subsection 1.3) with measured probabilities, thus silently invoking the Born rule.
@@
A solution of the measurement problem is indeed impossible unless one has a precise mathematical definition of when and how a mathematical model of a detector produces a definite response. This is the reason why, 100 years after the inception of quantum mechnaics in terms of states and observables, the quantum measurement problem is still unsolved.

### v2

@@ 2.3 Paradoxes and measurement context in Born’s rule
Already in 1926, Pauli [62, p.347] noticed the paradox that his objective position probability and Born’s objective momentum probability interpretation following from a spectral interpretation of the objective expectation form (BR-OE) of the Born rule are mutually incompatible. He expressed it in the following words: ”Man kann die Welt mit dem $p$-Auge und man kann sie mit dem $q$-Auge ansehen, aber wenn man beide Augen zugleich aufmachen will, dann wird man irre”
The 1927 paper by Jordan [37, p.811] cites Pauli [61] and extends Born’s second assumption further to an objective, measurement independent probability interpretation of inner products (probability amplitudes) of eigenstates of two arbitrary operators in place of position and momentum. without being aware of the conceptual problem this objective view poses when applied to noncommuting operators.
@@
2.4 Collapse (state reduction)
Tied to the measurement formulations of the Born rule is a controversial collapse (or state reduction) postulate about the state after a measurement. State reduction was introduced informally in Heisenberg [31, p.186] (”Jede Ortsbestimmung reduziert also das Wellenpaket wieder auf seine ursprüngliehe Grösse”) and formally established by the authority of Dirac in several editions of his book:
”The state of the system after the observation must be an eigenstate of [the observable] $\alpha$, since the result of a measurement of $\alpha$ for this state must be a certainty.” (Dirac [21, p.49], first edition, 1930)
”Thus after the first measurement has been made, the system is in an eigenstate of the dynamical variable $\xi$, the eigenvalue it belongs to being equal to the result of the first measurement. This conclusion must still hold if the second measurement is not actually made. In this way we see that a measurement always causes the system to jump into an eigenstate of the dynamical variable that is being measured, the eigenvalue this eigenstate belongs to being equal to the result of the measurement.” (Dirac [22, p.36], third edition, 1936)
Dirac’s statement was discredited (without much success) in 1958 by another authority, Landau & Lifshitz [42, Section 7], who remark that the state after the measurement is in general not an eigenstate, This is apparent in the modern notion of a quantum process (see, e.g., Chuang & Nielsen [16] or Mohseni et al. [48]), which gives the correct post-measurement state for arbitrary POVM measurements.
@@ 3 The Born rule today
For example, Schonfeld [66] assumes in his elementary (but somewhat heuristic) ’derivation’ of Born’s rule on p.4 the seemingly self-evident fact that the thermal randomness of subcritical droplet formation ”exists independent of any measurement axioms, i.e., independent of whether anyone observes it.” While his description gives some insight into the origin of the Born rule, the problem of deriving how random nucleation (rather than a mixture of all possible mucleations) arises from a quantum statistical model (cf. Subsection 3.5) is not even touched. The traditional statistical account of nucleation by Langer [41] is purely classical. A quantum derivation of nucleation was attempted, e.g., by Lombardo et al. [44]. But (like many others) they assume without discussion that (under reasonable conditions) the ”Wigner function can then be interpreted as a classical probability distribution for coordinates and momenta”, an assertion that can be maintained as referring to a statistical distribution only by invoking the condensed version (BR-C) of the Born rule.
In order not to follow into the trap of circularity, it is necessary to make the place explicit where measurement enters, and to define the properties a detector must have in order that its observations count as measurements. Therefore we begin with a precise conceptual framework within which some versions of the Born rule can be derived.
Following Neumaier & Westra [54, Part II], we now give precise definitions for quantum expectations, statistical expectations, and quantum measurement devices, and derive from it the formulations (BR-POVM) and (BR-C) of the Born rule.
@@ 3.1 The detector response principle
As shown in Neumaier & Westra [54, Part II], the DRP leads naturally to all basic concepts and properties of modern quantum mechanics. In particular, it gives a precise operational meaning to quantum states, quantum detectors, quantum processes, and quantum instruments. This gives a perspective on the foundations of quantum mechanics that is quite different from the well-trodden path followed by most quantum mechanics textbooks.
@@ 3.2 The statistical interpretation of quantum mechanics
The key result for the theory of quantum measurements is the following detector response theorem, proved in [54, Section 4.1.2].
@@ 3.3 Derivation of the Born rule
It is shown in [54, Part II] that the DRP together with techniques from quantum tomography allow one to reconstruct from the DRP the complete edifice of quantum mechanics, including its dynamical and spectral consequences. In particular, the DRP is precisely what is needed to obtain the concepts and results from quantum information thoery.
@@ 3.4.4 Domain of validity of the universal form
$\bullet$ As first observed by Heisenberg [32, p.25], the Born rule implies a tiny but positive probability that an electron bound to an atom is detected light years away from the atom: ”Das Resultat ist aber merkwürdiger, als es im ersten Augenblick den Anschein hat. Bekanntlich nimmt $\psi^{*}\psi$ exponentiell mit wachsendem Abstand vom Atomkern ab. Also besteht immer noch eine endliche Wahrscheinlichkeit dafür, das Elektron in sehr weitem Abstand vom Atomkern zu finden.” Therefore $|\psi(x)|^{2}$ is unlikely to be the exact probability density for being detected at $x$, as (BR-US) would require. This indicates that for observables with continuous spectrum, (BR-US) must be an idealization.
@@ 3.4.5 Domain of validity of the condensed form
The statistical interpretation based on the Born rule in the condensed form (BR-C) and in the POVM form (BR-POVM) are derivable from the DRP, and give a correct account of the actual experimental situation. They apply without idealization to results of arbitrary quantum measurements in the sense of the DRP. For a large number of examples see Neumaier & Westra [54, Section 6.2].
However, (BR-C) and (BR-POVM) still have the same limitations as the DRP, and cannot be applied to measurement of the kind discussed at the beginning of this subsection, such as macroscopic measurements or the measurements of spectral lines. For example, spectroscopy-based high precision measurements such as the gyromagnetic ratio of the electron are not quantum measurements but measurements of numerical parameters in specific quantum models; see the detailed discussion in Neumaier & Westra [54, Section 5.2].
@@ 3.5 What is missing in the foundations?
At every steadily progressing now, Nature somehow produces from what has been up to now (defined by what is observable, at least in principle) that what has been up to a while later.
Physics is supposed to quantitatively describe and understand this in terms of physical theories and models, with which predictions can be made, and in terms of experiments that evaluate some of the details of what has been up to now, with which such predictions can be prompted, verified, or refuted. But Nature proceeds without caring about our physical models and the knowledge built into these, and has done so long before the first living observers existed.
The fundamental physical state $\lambda(t)$ is a complete description of the (to us unknown) collection of everything that has been up to time $t$. The $\lambda(t)$ for all $t$ fully describe what is. This defines an effective ontology based on what is observable in principle.
The (to us unknown) fundamental dynamical law of physics determines how $\lambda(t)$ for $t>t_{0}$ is related to $\lambda(t_{0})$.
@@
Curiously, at present, our physical models are not in terms of what is (i.e., of approximations to $\lambda(t)$, part of which is observed) but in terms of states $\psi$ or $\rho$ and observables $X$ only indirectly related to what is. This indirect relation between the quantum theory of states and observables and reality is is given by one of the versions of the Born rule.
The above notion of what is must be distingushed from the theoretical ontology of certain interpretations of quantum mechanics: In the Copenhagen interpretation, what is is the classical world, delineated from the quantum world by a movable Heisenberg cut. In QBism (Fuchs [27, Section 3]), what is (in the present sense) is degraded to the world of a metaphysical society of agents governed by moral imperatives of how they should bet. In many world interpretations, a preferred ontological status is given to a postulated universal wave function, a theoretical figment of abstraction, whereas what is (in the present sense) is degraded to a particular world amidst infinitely many worlds, namely the single distinguished world that is our actual past, and from which we have to learn everything we can know about Nature, independent of any weight one might give to this world. In Bohmian mechanics, infinitely precise positions and the exact wave function are declared to have a preferred ontological status – a poor, theoretical substitute for the colorful what is that we experience. In each case, nothing real (in the sense of our what is) is in the quantum theory (of states and observables) itself, a state of affairs criticized by Einstein throughout his life.
@@
The Born rule depends on a notion of measurement that has no formal counterpart in quantum mechanical models. This means that we are currently unable to explain theoretically which formal quantum models of pieces of matter describe a quantum detector. The theoretical explanation for why certain objects are quantum detectors constitutes the so-called quantum measurement problem.
@@
If we accept the DRP, the problem is to give mathematical conditions on certain individual quantum objects described by microscopic quantum mechanics that guarantee that the DRP holds for a suitable mathemetcal definition of what it means for the quantum model to respond at a particular time. This is a nontrivial unsolved problem in the statistical mechanics of detector properties.
In 1962, Daneri et al. [17] initiated the discuss of quantum detectors by means of quantum statistical mechanics. However, they assume the objective scattering form (BR-Sc) of the Born rule: ”the intermediate state (depending on the direction in which the particle is scattered) is the state in which ions are created inside the $k$-th counter” (p.308). In addition, they assume that the detector is dissipative (p.307). The latter (but not some version of the Born rule) could be avoided by using a metastable quantum device coupled to a heat bath with a continuous frequency spectrum, along the lines of Caldeira & Leggett [15].
The work by Allahverdyan et al. [2] goes some way towards resolving the problem. But they have to equate at some points the formal probabilities of statistical mechanics (cf. Subsection 1.3) with measured probabilities, thus silently invoking the Born rule. This is done at the beginning of Section 11, where they say that ”the statistical interpretation of quantum mechanics emphasizes the idea that this theory, whether it deals with pure or mixed states, does not govern individual systems but statistical ensembles of systems”. But it describes statistical ensembles only if the formal probabilities are taken to be observed frequentist probabilities, which requires Born’s rule. Thus the derivation is circular. The problem is still present in their recent paper [3, p.254], where we read: ”we acknowledge that each individual run of the measurement produces a single macroscopic indication of the pointer of the apparatus which may be read or registered. This statement, although intuitive, is not a consequence of the mere quantum principles and has a macroscopic nature. We will accept its inclusion in quantum theory as a postulate (motivated by the macroscopic size of M)”. In fact, when they write that ”what we call ’quantum state’, whether pure or not, is in general an abstract mathematical object which allows us to make probabilistic predictions about the outcomes of any experiment”, they assume much more, namely the condensed form (BR-C) of Born’s rule! In effect they show that (BR-C) assumed for quantum statistical mechanics implies the Born rule for individual events.
The decoherent history approach of Gell-Mann & Hartle [28] for the derivation of classical probability suffers from the problem of having to reinterpret formal smeared Wigner functions as classical probability densities (cf. my introduction to Section 3), whose statistical interpretation needs (BR-C). Indeed, they assert on p.3378 that something is missing: ”we may begin to tackle the deep problem of introducing individuality into quantum mechanics. Actual alternative histories deal, of course, in large part with individual objects such as our galaxy, the Sun, the Earth, biological organisms on the Earth, and so forth. Yet discussions of quantum mechanics up to now have typically treated such individual objects only as external systems, labeled as ’observers’ and ’pieces of apparatus’.”
The main problem behind these difficulties is the need to find a rigorous derivation of a dissipative classical dynamics for a single pointer (or switch) without a silent recourse to the Born rule. A customary way to proceed (following Möhring & Smilansky [47]) is to start with the influence functional approach of Feynman & Vernon [24]. However, this approach has the Born rule built in into the derivation (p.122) of the basic functional integral representation of the transition probabilities.
@@
A solution of the measurement problem is indeed impossible unless one has a precise mathematical definition of when and how a mathematical model of a single quantum detector produces a definite response.666 ”There is no consistent quantum measurement theory, unless an appropriate reinterpretation of what it means for an observable to have a definite value can be found.” (Busch & Lahti [14, p.375]) This is the reason why, 100 years after the inception of quantum mechanics in terms of states and observables, the quantum measurement problem is still unsolved. It is the missing piece in the foundations of quantum mechanics.